Invert the matrix 123 and show that the 645 1798 many times Its Yonverse is the identity matry..
Answer #1

let A be A the matrix [1 2 3 6 4 5 7 9 8 & be the matrix of let of cofactors A. [-13 In 1-2 – 13 – 13 13 267 5 -8] det = - 13As A | - و3%- % % 35- و173 3-37 2 = 3- ا - 5 4 | . د له و ا + ا رابر) ا ماه | -ابرا | او در اد م د م ن و ( اله ماابي | 41 + )